ITES2001 Fourth Italian-Spanish conference on General Topology and its applications Bressanone, 27-30 June 2001

Embeddings into Pow ℕ/ Fin and extension of automorphisms

A. Bella, A. Dow and P. Ursino Università di Catania, Italy

Given a Boolean algebra 𝔹 and an embedding e : 𝔹 ⟶ Pow ℕ/ Fin, we consider the possibility of extending each or some automorphism of 𝔹 to the whole Pow ℕ/ Fin. Among other things, we show that there are embeddings for which no non-trivial automorphism can be extented.